A sample of helium saturated with water vapour exerts a total pressure of 103.17 kPa. If the pressure of the helium is 99.39 kPa, what is the temperature?
1
Expert's answer
2020-06-19T07:01:34-0400

At equillibrium temperature of water vapour is 100°c

So we have,, P1/T1=P2/T2

=103170/100=99390/T2

=T2=99390/1031.7

=96.3 degrees.